OpenStudy (anonymous):

Mona has a triangular flag. She wants to cut out a similar flag that has dimensions one-fourth of the original flag. If the area of the original flag is 320 square inches, what is the area of the new flag that Mona cuts out? Answer 16 in2 20 in2 32 in2 80 in2

OpenStudy (anonymous):

320 / 4 = 80

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks

Directrix (directrix):

If two polygons are similar, the square of their scale factor is equal to the ratio of their areas. (1/4)^2 = x / 320 where x is the area of the smaller flag 1/16 = x / 320 x = 20